OpenStudy (anonymous):

let V1 = (2,-6) and V2 = (-4,7) compute the unit vectors in the direction of absolute value V1 and absolute value V2.

OpenStudy (anonymous):

Do you know how to calculate a unit vector?

OpenStudy (anonymous):

\[unit vector v=\frac{ vector v }{ \left| v \right| }\]

OpenStudy (anonymous):

\[v1=\frac{ 2i-6j }{\sqrt{2^{2}+\left( -6 \right)^{2}} }\]

OpenStudy (anonymous):

calculate it similarly you can find unit vector V2